# MORPHeus: a Multimodal One-armed Robot-assisted Peeling System with Human Users In-the-loop

### v1

@@ I Introduction
Conventionally, peeling is considered a bimanual task in which one arm stabilizes the food item while the other arm executes the peeling action [19]. Most real-world robotic caregiving systems [20] use one robot arm mounted on the wheelchair because mounting two robot arms on wheelchairs is practically challenging due to ergonomic and transfer challenges for care recipients [21]. We use an assistive cutting board inspired by how individuals with one functional hand prepare meals. An assistive cutting board [22], also known as an adaptive cutting board, is a board with multiple fixtures that can secure a food item (shown in Fig 1), so the user can peel with one arm. In addition, when users peel, they might have different preferences. For example, a user might want a half-peeled cucumber if they want to prepare cucumber bites with some texture provided by the skin. We want to ensure that their preferences about peeling coverage (unpeeled, half-peeled, top-peeled, etc.) are taken into account.
@@
Our key insight is that multimodal perception with human-in-the-loop long-horizon planning and compliant control can enable a one-armed robot to perform the peeling task for a wide variety of food items while taking users’ preferences into consideration. Using this insight, we develop a system that includes: (1) a multimodal active perception module using visual, haptic, and vibration sensing modalities to estimate the peeling state of a food item during task execution; (2) a Large Language Model (LLM)-based interface to convert a user’s natural language commands to long-horizon plans with a human user in-the-loop; (3) a compliant controller for executing the peeling motion and adapting to a wide variety of food items with varying physical properties. Check out the MORPHeus system at https://emprise.cs.cornell.edu/morpheus/.
@@
Figure 2: Overview of the building blocks of MORPHeus: Multimodal perception (Sec. III-A) module uses visual, vibration, and haptic inputs to determine whether the food is peeled, Human-in-the-loop Long-horizon Planning (Sec. III-B) enables the robot to perform task planning while catering to a user’s preference and perform anomaly detection using user’s feedback, and Compliant Control (Sec. III-C) to execute peeling actions for various food items.
@@
  • A multimodal active perception model incorporating visual, haptic, and vibration inputs for peeling.
@@
  • A dataset of human peeling actions on a wide variety of food items and an open-hardware design of a peeler with sensors for visual, haptic, and vibration modalities.
@@ II-A Meal Preparation with Robots
Meal preparation can be broken down into multiple substeps, including cutting [12, 13], peeling [19], and cooking [16]. Xu et al. [12] developed a cutting system to cut food with solid cores, such as avocado. Liu et al. [16] proposed a model that learned stir-fry motions from human demonstrations. These works either assume that the food items have already been peeled or ignore the peeling problem. There is also research on peeling bananas with a dexterous hand [23], peeling grapes with a surgical robot [24], and peeling a grapefruit with a knife [25]. Watanabe et al. [26] and Dong et al. [19] used peelers for peeling, similar to our setup. However, they both used a dual-arm setup: one arm stabilizes the food items, and the other peels. Watanabe et al. [26] used 2D visual perception to determine whether the food was peeled and tested their system with a single food item. Dong et al. [19] added a point cloud to obtain 3D information and expanded the scope to 5 food items. Our work uses visual, haptic, and vibration modalities to determine whether the food item is peeled, takes a variety of assistive cutting boards into consideration, and performs experiments on 12 vegetables and fruits that cover a wide range of physical properties, with only one robot arm. Additionally, our work differs from previous work in the fact that we perform the peeling task based on user preferences about the peeling coverage.
@@ III-A1 Perception Model
Figure 3: Model structure for multimodal perception: The network uses image, vibration, and haptics as input, and gives a binary output for peeled/unpeeled classification ($O$).
@@ III-B Human-in-the-loop Long-horizon Planning
Prompt Specification: We provide GPT-4 with a structured prompt as follows. The prompt describes (a) the role of GPT, (b) the robot actions, (c) the environment, (d) the format of the output, (e) example of a customized prompt with the same structure with the generated domain PDDL, and (f) user inputs, such as peeling goals and feedback. We show an example of the prompt inside the gray box III-B-1. The prompt with these components gives GPT the context to reason for this specific peeling task, and also constraints its output syntax by giving examples of PDDL files. Prior research, such as  [50, 51], has highlighted the proficiency of Large Language Models (LLMs) in utilizing human feedback to detect and correct anomalies in generated content. In our planning approach, we implement an anomaly detection with user’s feedback module to address potential execution errors. At critical junctures in the plan, this module will seek human feedback to detect or address potential anomalies. After completing the peeling of a certain area or a flip, the process pauses to ask the user some questions regarding whether the action has succeeded. For example, the robot may ask “Has the rotation action succeeded?” If the user identifies an anomaly, the LLM engages the user with a series of binary questions, such as “Is the food item between the jaws?”, to update its understanding of the environment. Following this, the LLM replans to rectify the previous failure before proceeding, assume undo is possible.
@@ IV-C Compliant Control
[…]

### v2

@@ I Introduction
Conventionally, peeling is considered a bimanual task in which one arm stabilizes the food item while the other arm executes the peeling action [19]. Most real-world robotic caregiving systems [20] use one robot arm mounted on the wheelchair because mounting two robot arms on wheelchairs is practically challenging due to ergonomic and transfer challenges for care recipients [21]. We use an assistive cutting board inspired by how individuals with one functional hand prepare meals. An assistive cutting board [22], also known as an adaptive cutting board, is a board with multiple fixtures that can secure a food item (shown in Fig 1), so the user can peel with one arm. In addition, when users peel, they might have different preferences. For example, a user might want a half-peeled cucumber if they want to prepare cucumber bites with some texture provided by the skin. We want to ensure that their preferences about peeling coverage (fully-peeled, half-peeled, top-peeled, etc.) are taken into account.
@@
Our key insight is that multimodal perception with human-in-the-loop long-horizon planning and compliant control can enable a one-armed robot to perform the peeling task for a wide variety of food items while taking users’ preferences into consideration. Using this insight, we develop a system that includes: (1) a multimodal active perception module using visual, force, and vibration sensing modalities to estimate the peeling state of a food item during task execution; (2) a Large Language Model (LLM)-based interface to convert a user’s natural language commands to long-horizon plans with a human user in-the-loop; (3) a compliant controller for executing the peeling motion and adapting to a wide variety of food items with varying physical properties. Check out the MORPHeus system at https://emprise.cs.cornell.edu/morpheus/.
@@
Figure 2: Overview of the building blocks of MORPHeus: Multimodal perception (Sec. III-A) module uses visual, vibration, and force inputs to determine whether the food is peeled, Human-in-the-loop Long-horizon Planning (Sec. III-B) enables the robot to perform task planning while catering to a user’s preference and perform anomaly detection using user’s feedback, and Compliant Control (Sec. III-C) to execute peeling actions for various food items.
@@
  • A multimodal active perception model incorporating visual, force, and vibration inputs for peeling.
@@
  • A dataset of human peeling actions on a wide variety of food items and an open-hardware design of a peeler with sensors for visual, force, and vibration modalities.
@@ II-A Meal Preparation with Robots
Meal preparation can be broken down into multiple substeps, including cutting [12, 13], peeling [19], and cooking [16]. Xu et al. [12] developed a cutting system to cut food with solid cores, such as avocado. Liu et al. [16] proposed a model that learned stir-fry motions from human demonstrations. These works either assume that the food items have already been peeled or ignore the peeling problem. There is also research on peeling bananas with a dexterous hand [23], peeling grapes with a surgical robot [24], and peeling a grapefruit with a knife [25]. Watanabe et al. [26] and Dong et al. [19] used peelers for peeling, similar to our setup. However, they both used a dual-arm setup: one arm stabilizes the food items, and the other peels. Watanabe et al. [26] used 2D visual perception to determine whether the food was peeled and tested their system with a single food item. Dong et al. [19] added a point cloud to obtain 3D information and expanded the scope to 5 food items. Our work uses visual, force, and vibration modalities to determine whether the food item is peeled, takes a variety of assistive cutting boards into consideration, and performs experiments on 12 vegetables and fruits that cover a wide range of physical properties, with only one robot arm. Additionally, our work differs from previous work in the fact that we perform the peeling task based on user preferences about the peeling coverage.
@@ III-A1 Perception Model
Figure 3: Model structure for multimodal perception: The network uses image, vibration, and force as input, and gives a binary output for peeled/unpeeled classification ($O$).
@@ III-B Human-in-the-loop Long-horizon Planning
Prompt Specification: We provide GPT-4 with a structured prompt as follows. The prompt describes (a) the role of GPT, (b) the robot actions, (c) the environment, (d) the format of the output, (e) example of a customized prompt with the same structure with the generated domain PDDL, and (f) user inputs, such as peeling goals and feedback. We show an example of the prompt inside the gray box III-B-1. The prompt with these components gives GPT the context to reason for this specific peeling task, and also constraints its output syntax by giving examples of PDDL files. Prior research, such as  [50, 51], has highlighted the proficiency of Large Language Models (LLMs) in utilizing human feedback to detect and correct anomalies in generated content. In our planning approach, we implement an anomaly detection with user’s feedback module to address potential execution errors. At critical junctures in the plan, this module will seek human feedback to detect or address potential anomalies. After completing a rotation or a flip, the process pauses to ask the user some questions regarding whether the action has succeeded. For example, the robot may ask “Has the rotation action succeeded?” If the user identifies an anomaly, the LLM engages the user with a series of binary questions, such as “Is the food item between the jaws?”, to update its understanding of the environment. Following this, the LLM replans to rectify the previous failure before proceeding, assume undo is possible.
@@ IV-C Compliant Control
Start states NAG1 Avg. Action Max. Action Min. Action
GPT-Domain &\&& Fixed-Problem2